Consider a price-taking firm that has total fixed cost of $500 and faces a market-determined price of $10 per unit for its output. The wage rate is $175 per unit of labor, the only variable input. Using the following table, fill in the columns and answer the question below.

(1)

(2)

(3)

(4)

(5)

(6)

Units of labor

Output

Marginal product

Marginal revenue product

Marginal cost

Profit

1

10

2

30

3

60

4

100

5

150

6

190

7

210

8

225

9

235

10

240

The manager should produce _____ units of output to maximize profit.
Answer #1

Marginal product is the change in the total output due to employment of additional unit of resource i.e. in this problem an additional unit of labor.

Revenue is the product of price and quantity sold i.e. Revenue = Price ×  Quantity sold

Marginal revenue product is the change in the total revenue due to employment of an additional unit of input i.e. in this case an additional unit of labor.

Total cost = Fixed cost + variable cost

Variable cost = Units of labor × wage rate per labor

Profit = Revenue - total cost

Units of labor Output Marginal product Marginal revenue product Marginal cost Profit
1 10 _ _ _ -$575
2 30 20 $200 $175 -$550
3 60 30 $300 $175 -$425
4 100 40 $400 $175 -$200
5 150 50 $500 $175 $125
6 190 40 $400 $175 $350
7 210 20 $200 $175 $375
8 225 15 $150 $175 $350
9 235 10 $100 $175 $275
10 240 5 $50 $175 $150

The calculation related to the above table is shown as follows:-

Units of labor Output Marginal product Revenue Marginal revenue product Total cost Marginal cost Profit
1 10 10 × $10 = $100 _ $675 _ $100 - $675 = - $ 575
2 30 30-10 = 20 30 × $10 = $300 $300 - $100 = $200 $850 850 - 675 = 175 $300 - $850 = - $ 550
3 60 60 - 30 = 30 60 × $10 = $600 $600 - $300 = $300 $1,025 1025 - 850 = 175 $600 - $1025 = -$425
4 100 100 - 60 = 40 100 × $10 = $1000 $1000 - $600 = $400 $1,200 1200 - 1025 = 175 $1000 - $1200 = -$200
5 150 150 - 100 = 50 150 × $10 = $1500 $1500 - $1000 = $500 $1,375 1375 - 1200 = 175 $1500 - $1375 = $125
6 190 190 - 150 = 40 190 × $10 = $1900 $1900 - $1500 = $400 $1,550 1550 - 1375 = 175 $1900 - $1550 = $350
7 210 210 - 190 = 20 210 × $10 = $2100 $2100 - $1900 = $200 $1,725 1725 - 1550 = 175 $2100 - $1725 = $375
8 225 225 - 210 = 15 225 × $10 = $2250 $2250 - $2100 = $150 $1,900 1900 - 1725 = 175 $2250 - $1900 = $350
9 235 235 - 225 = 10 235 × $10 = $2350 $2350 - $2250 = $ 100 $2,075 2075 - 1900 = 175 $2350 - $2075 = $275
10 240 240 - 235 = 5 240 × $10 = $2400 $2400 - $2350 = $50 $2,250 2250 - 2075 = 175 $2400 - $2250 = $150

The manager should produce 210 units of output to maximize the profit